Using the following data($ billions) for a given year and assuming that GDP equals GNP,calculate (a) gross domestic product (b) national income (c) net national income (d) indirect business taxes.

1.consumption spending = $8000

2.interest income = 500

3.rental income = 200

4.investment spending = 2,500

5.profits = 2,000

6.wages salaries = $ 7000

7.depreciation =1550

8.government spending =2000

9.net export spending = -450


1
Expert's answer
2022-03-03T12:07:37-0500

If GDP is the same as GNP


A) "GDP=C+I+G+X_n"

"GDP=\$8000+\$2500+\$2000-\$450"

"GDP=\$12050"


B) National income=

"w+r+i+\pi+Tb"

Where Tb is indirect business tax

"NI=\$7000+\$200+\$500+\$2000+Tb"TbTbTbTbTbTbTbTbTTbTbTbTbTbTbTTbTbTbTb

"NI=\$11700+Tb"


C) Net national product= "GDP-D"

Where D is depreciation.

"NNP=\$12050-\$1550"

"NNP=\$10500"


D) Indirect business tax

"NI=NNP-Tb"

but "NI=\$11700+Tb"


"\therefore \$11700+Tb=\$10500-Tb"

"Tb=-\$600"


"NI=\$11700-\$600=\$11100"